Solve the following system of linear equations. (If the system has no solution, enter NONE for each answer. If the system is dependent, enter INFINITELY MANY for each answer.)
x-y=4
3x-3y=12
x = ?
y = ?
Pls show me the steps on how to get the answer. thanks

Answer by stanbon(75887)

x-y=4
3x-3y=12
----
Divide thru the 2nd equation by 3:
You get:
x-y = 4
x-y = 4
----
The two equations are the same equation
Solve for "y":
y = x-4
---
The solution is all points of the form: (x,x-4)
where x is a Real Number.
===================================
Cheers,
Stan H.
===============